The Probability that a first-time tourist to the city of chicago will visit the Art Institute is 0.4, will visit the museum of science and industry is 0.3, and will visit both is 0.1 assume a first-tiem tourist to chicago is randomly selected. Find the probaability that the tourist will visit the art institute or museum of science and industry?
1
Expert's answer
2011-10-27T08:25:30-0400
Let’s saythat P(A) = 0.4 – probability of visiting the Art InstituteP(M )=0.3 – probability of visiting the museumof science and industry

P(A and M)probability of visiting both of them

P(A or M)- probability of visiting the art institute or museum ofscience and industry
We’ll use aformula: P(A or M)=P(A)+P(M)-P(A and M)

Let’s put givenvalues into formula: P(A or M)=0.4+0.3-0.1=0.6



Answer 0.6